From the top of the tower of height 400 m, a ball is dropped by a man. Simultaneously from the base of the tower, another ball is thrown up with velocity 50 m/s. At what distance will they meet from the base of the tower ?

Grade:12

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
Let us assume that the balls meet at a distance of x m from the base of the tower.
Hence the velocity achieved by the ball by the time it reaches the distance (400-x) from the top is:
v^2=u^2+2gh\Rightarrow v^2=0^2+2\times9.8\(400-x\)\Rightarrow v=\sqrt 19.6\(400-x\)Lets assume that it takes t sec to cover this distance, hence, we have:
400-x=0\times t+\frac{1}{2}gt^2
Similarly for the ball thrown from the bottom, we have:
x=50t-\frac{1}{2}gt^2
so, we get:
400=50t\Rightarrow t=8 sec
hence,
x=50\times 8-\frac{1}{2}\times 9.8\times 8^2=400-313.6=86.4 m
Regards